A college conducts both face to face and online classes that are intended to be identical. Given below are
the distributions of the two groups in terms of marks scored in a statistics examination.
Face to face Class 0 – 10 10 – 20 20 – 30 30 – 40 40 – 50 50 – 60 60 – 70
No. of Students 5 10 20 25 15 15 10
On-line Class 0 – 10 10 – 20 20 – 30 30 – 40 40 – 50 50 – 60 60 – 70
No. of Students 10 25 30 40 30 20 15
Test the hypothesis that the mean marks are statistically equal at the 5% level of significance and write a 5%
confidence interval estimate of the difference
1
Expert's answer
2020-10-28T17:04:21-0400

Solution:









Face to Face Class

"Mean, \mu_1 = {\sum f1x \over \sum f1}""= {3700 \over 100} = 37"

"var , \sigma_1 ^2 = {\sum f1{(x- \mu)}^2 \over \sum f1}""={26600 \over 100}= 266"

On-line Class

"=Mean, \mu_2 = {\sum f2x \over \sum f2}""= {6000 \over 170} = 35.2941"

"var , \sigma_2 ^2 = {\sum f2{(x- \mu)}^2 \over \sum f2}""= {46485.29412 \over 170} = 273.4429"

Hypothesis Testing.

"H_0 : \mu_1 = \mu_2" vs "H_1: \mu_1 \not = \mu_2" at 5% level of significance


Test statistic:


"= {\mu_1 - \mu_2 \over \sqrt{{\sigma_1^2 \over n_1} + {\sigma_2^2 \over n_2}}}"

"= {37 -35.2941 \over \sqrt{{266 \over 100}+ {273.4429 \over 170}}} = 0.8257"

Z0.025=1.96


Conclusion:

Since the calculated value of Z is less than the table value of Z, we fail to reject the null hypothesis. Therefore, the mean marks are statistically equal.


Confidence Interval.

"=(\mu_1 - \mu_2) \pm Z_{\alpha \over2}{\sqrt{{\sigma_1^2 \over n_1} + {\sigma_2^2 \over n_2}}}"

"=(37-35.2941) \pm 1.96 \sqrt{{266 \over 100}+{273.4429 \over 170}}"

"= 1.7059 \pm 4.0494"

Ans: "-2.3435 \le \mu_1 -\mu_2 \le 5.7553"